decide whether the pairof lines is parallel,perpendicular, or neither
5x+2y=4
2x-5y=5

Answer by mananth(16949)

5x+2y=4
2y=4-5x
divide by 2
2y/2=4/2-5x/2
y=-5x/2 +2
The slope of this line = -5/2
2x-5y=5
-5y=5-2x
divide by -5
-5y/-5=5/-5 -2x/-5
y=-1+2x/5
y=2x/5-1
the slope of this line = 2/5
slope of od * slope of other = -1
m1*m2=-1
so they are perpendicular.